OpenStudy (anonymous):

Greatest Common Factor help? 1) 12x and 26 GCF; 2)18 x^5 and 15 x^5 GCF; 3)12 a^5b and 40 a^5b^5 GCF; 4).15 f^2 g h^3 and 60 f^2 h^3

jhonyy9 (jhonyy9):

1). 12x = 2*2*3*x 26 = 2*13 ------------------ GCF = ? how do you think what will be ?

jhonyy9 (jhonyy9):

common Jessie ,please cooperate anything

jhonyy9 (jhonyy9):

what number is common there in these two numbers ?

OpenStudy (anonymous):

So its 2

jhonyy9 (jhonyy9):

yes sure is right

jhonyy9 (jhonyy9):

was easy ,yes ?

jhonyy9 (jhonyy9):

so after this eaxmple what is your first exercise how do you think what will be for 2, ?

OpenStudy (anonymous):

I don't understand how to get the factors with the exponents on them

jhonyy9 (jhonyy9):

in case of second exercise will be easy because these exponents are the same

OpenStudy (anonymous):

yeah but how do you find the factors? That's what I don't understand

jhonyy9 (jhonyy9):

so looke it if there is x^3 the factors will be x*x*x

jhonyy9 (jhonyy9):

for 18 = 2*3*3

jhonyy9 (jhonyy9):

for 15 = ?

jhonyy9 (jhonyy9):

how can you getting the 15 like product of what numbers ?

jhonyy9 (jhonyy9):

divide 15 by 3 what will get ?

jhonyy9 (jhonyy9):

ok i will make it again but the 3rd will calculi it you ,ok ?

jhonyy9 (jhonyy9):

18x^5 = 2*3*3*x*x*x*x*x 15x^5 = 2*2*3*x*x*x*x*x --------------------------- GCF = ?

jhonyy9 (jhonyy9):

what will be common there in these two ,,lines" ?

jhonyy9 (jhonyy9):

so 2 is common ?

OpenStudy (anonymous):

Thanks for the help but I figured them out with my brother

jhonyy9 (jhonyy9):

18x^5 = 2*3*3*x*x*x*x*x 15x^5 = 5*3*x*x*x*x*x ---------------------------- GCF = 2 is common ? not is 3 is common ? yes but just one time 5 is common ? not is x is common ? yes but the five time ------------------------------- so than the GCF = 3*x*x*x*x*x = 3x^5

jhonyy9 (jhonyy9):

now this is clearly

jhonyy9 (jhonyy9):

hope so much that was these on your help good luck bye